what is x?
-2/10x + 2/5(x - 2)= -1

Answer by rfer(16322)

-2/10x+4/10x-2/5=-1
1/5x-2/5=-1
1/5x=-1+2/5
1/5x=-3/5
x=-3/5*5/1
x=-15/5
x=-5